# A flexible framework for structural plasticity in GPU-accelerated sparse spiking neural networks

### v1

@@ 2.1.3 Synaptic rewiring for topographic maps
Figure 3: Illustration of data structures used for one custom connectivity update according to the synaptic rewiring model for topographic maps. Each panel shows a network sketch, the ragged matrix (number of presynaptic neurons $\times$ row length), the bitfield (number of presynaptic neurons $\times$ number of postsynaptic neurons), and pseudocode. (A) Initial state. Five presynaptic neurons, sources $s_{j}$ with $0\leq j<5,$ are connected to four postsynaptic neurons, targets $t_{i}$ with $0\leq i<4$, using sparse connectivity with a maximum row length of three. Weights below the threshold $g_{\theta}$ are indicated in light red, others in light blue. A synapse slot from $s_{j}$ to $t_{i}$ is here referred to as $\left(s_{j},t_{i}\right)$. The bitfield is cleared. (B) Host assigns six rewiring attempts randomly to three presynaptic neurons ($s_{0}$, $s_{2}$, $s_{4}$). The presynaptic neurons then in parallel select their postsynaptic targets randomly and set the corresponding bits as candidates for rewiring attempts in the bitfield. Four existing synapses are selected for elimination (green; $\left(s_{0},t_{0}\right)$, $\left(s_{0},t_{2}\right)$, $\left(s_{4},t_{2}\right)$, $\left(s_{4},t_{3}\right)$) and two empty synapse slots for formation (yellow; $\left(s_{2},t_{2}\right),\left(s_{4},t_{0}\right)$). (C) Elimination attempts. $\left(s_{0},t_{0}\right)$ has a low elimination probability because of the strong weight (light blue) and remains, the three others have a higher elimination probability because of the lower weights (light red); $\left(s_{0},t_{2}\right)$ gets eliminated reducing the row length of $s_{0}$ in the ragged matrix from two to one; $\left(s_{4},t_{2}\right)$ gets eliminated, and the emerging gap in the row of $s_{4}$ is filled by inserting the last element $t_{3}$; $\left(s_{4},t_{3}\right)$ remains. The elimination bits get cleared in the bitfield. (D) Formation attempts. Assuming a spatial proximity of $s_{2}$ and $t_{2}$ and therefore a high formation probability, a new synapse (purple) is formed at $\left(s_{2},t_{2}\right)$, increasing the row length of $s_{2}$ from zero to one, but not at $\left(s_{4},t_{0}\right)$ because $s_{4}$ and $t_{0}$ are farther apart. The formation bits get cleared in the bitfield.
@@
We develop a GeNN-compatible variation of the synaptic rewiring model proposed by [22], drawing inspiration from an implementation on the SpiNNaker platform [7]. This stochastic model combines a distance-dependent formation rule with a weight-dependent removal rule. The simulation loop performs custom connectivity updates at regular intervals according to a rewiring frequency. Each update triggers a fixed number of rewiring attempts between neurons from the presynaptic and postsynaptic population. While previous implementations [22, 7] handle each rewiring attempt separately, our algorithm aims to increase efficiency by parallelizing multiple attempts. figure 3 illustrates the hybrid CPU (host) and GPU (device) algorithm processing these attempts and highlights the data structures. First, serial host code (host_update_code) uniformly distributes the attempts across all neurons in the presynaptic population. As we sample with replacement, multiple attempts per presynaptic neuron are possible. In contrast to earlier implementations [22, 7], our version does not impose a fixed maximum in-degree (called synaptic capacity). GeNN represents connectivity in 2D-array structures; the rows correspond to presynaptic neurons and can be operated on independently and hence in parallel. Here, we introduce a bitfield as an auxiliary structure in a similar format to mark synaptic slots between specific pre- and postsynaptic neurons for wiring attempts. On the GPU (row_update_code), each presynaptic neuron accesses a unique section of the bitfield with as many bits as there are neurons in the postsynaptic population. The bits therefore encode synapse slots for potential synapses between the presynaptic neuron of the row considered and all postsynaptic neurons; each bit refers to a postsynaptic neuron index. The bitfield section is initialized to zero and then as many bits are set randomly according to a uniform distribution without replacement as there are wiring attempts assigned. Each synaptic slot is therefore considered at most once, unlike the earlier versions [22, 7] where each rewiring attempt is completely independent of the others.
@@
When the presynaptic neuron has set its bitfield row, the elimination rule is applied: A loop over all existing synapses of that presynaptic neuron identifies those that are set in the bitfield section, removes them dependent on an elimination probability compared to a random number and clears the respective bit. Weak synapses are more likely to be removed: the elimination probability is higher for depressed synapses with weights below the threshold than for strong, potentiated synapses. After all elimination attempts, the remaining set bits in the bitfield row point to synapse slots that were initially empty. Applying the formation rule, we then iterate over the whole bitfield row: If a bit is still set, we attempt to establish a synapse according to a Gaussian probability depending on the distance between the neurons and also clear that bit. The shorter the lateral distance between pre- and postsynaptic neuron, the more likely is a new connection. Our interpretation of the rewiring model does not enable the formation of multapses (i.e., multiple synapses per neuron pair [1]). If a synapse exists, only the elimination rule is applied. To sum up, each custom connectivity update initiates a fixed number of rewiring attempts between pre- and postsynaptic neuron populations. The outcome of each attempt is a newly formed connection, a removed connection, or no change to the connectivity.
@@ 2.2.1 Recurrent classifier
Our previous work [16, 17, 18] shows that recurrent SNN classifiers can be trained on datasets such as the Spiking Heidelberg Digits [33] and DVS gesture [21].
Here we train SNN classifiers with a single recurrently connected hidden layers of 512 adaptive LIF (ALIF) neurons on the DVS gestures dataset. Following the approach outlined by [34], we spatially downsample the dataset examples from $128\times 128$ to $32\times 32$ resolution and temporally downsample so that there is only one spike in each $1\,\mathrm{ms}$ simulation timestep.
In discrete time form, the membrane voltage ($v$), adaptation variable ($a$) and spiking output ($z$) of each hidden neuron in these models can be described by:
@@
where $\alpha=e^{-\frac{\textrm{dt}}{\tau_{\textrm{mem}}}}$represents the decay of the membrane with time constant $\tau_{\textrm{mem}}=20\,\mathrm{ms}$, $W^{\textrm{rec}}$ are the recurrent weights, $W^{\textrm{in}}$ are the weights connecting the input to the recurrent layer, $x^{t}$ are input spikes from the dataset, $v_{\textrm{thr}}=0.6$ is the baseline spiking threshold, $\rho=e^{-\frac{\textrm{dt}}{\tau_{\textrm{adapt}}}}$represents the adaptation variable decay with time constant $\tau_{\textrm{adapt}}=2000\,\mathrm{ms}$ and $\beta=0.0174$ is the adaptation strength. When neurons spike, $v_{\textrm{thr}}$ is subtracted from their membrane voltage $v$. In our classifiers, $W^{\textrm{rec}}$ and $W^{\textrm{in}}$ are sparse and initialised with a fixed probability of connection between each pair of pre and postsynaptic neurons e.g. we use $5\,\mathrm{\%}$ connectivity to indicate that there is a $5\,\mathrm{\%}$ chance of a pair of pre and postsynaptic neurons being connected.
@@
We calculate weight gradients $\Delta W^{\textrm{rec}}$ and $\Delta W^{\textrm{in}}$ using eligibility traces $\epsilon$ and $e$ calculated using e-prop [19] (shown here for $\Delta W^{\textrm{rec}}$) :
@@
where $\psi_{j}^{t}=\frac{0.3}{v_{\textrm{thr}}}\text{max}\left(0,1-\left|\frac{v_{j}^{t}-(v_{\textrm{thr}}+\beta a_{j}^{t})}{v_{\textrm{thr}}}\right|\right)$ is a surrogate gradient function, $\bar{z}^{t}$ is a version of $z^{t}$ low-pass filtered with a time constant of $\tau_{\textrm{mem}}$, $\eta$ is a learning rate, $B=\left(W^{\textrm{out}}\right)^{T}$, $\pi_{k}^{t}=\textrm{softmax}(y_{1}^{t},\ldots,y_{K}^{t})$, $\pi^{*,t}$ are one-hot encoded labels and $\bar{e}^{t}$ is a version of $e^{t}$ low-pass filtered with a time constant of $\tau_{\textrm{mem}}$.
@@
All gradients are applied to the underlying variables using the Adam optimizer [35] and $\Delta W^{\textrm{rec}}$ and $\Delta W^{\textrm{in}}$provide the inputs to the DEEP R algorithm described in section 2.1.2.
@@ B Neuron, input and plasticity models
Table 1: Topographic map model description and parameters following the guidelines of [36].
@@ F Synaptic plasticity
$B$ $1.2$ Ratio of potentiation to depression es defined in [37]
@@ G Structural plasticity
Elimination probability for depressed synapses ($g_{\text{syn}}<g_{\theta}$); value from [22] increased by factor
@@
Elimination probability for potentiated synapses ($g_{\text{syn}}\geq g_{\theta}$); value from [22] increased by factor
@@ H Simulation
The network model illustrated in figure 4A and its parameters closely match the earlier accounts on topographic map formation of [22, 7]. In this section, we provide a summary and emphasize the differences, while tables 1 and 2 give a complete overview of the model description and parameters used in our work in the format frequently used in computational neuroscience [36]. The model consists of two square layers with nodes located at grid positions: model neurons in the target layer and spike-generating devices in the source layer. Feed-forward connections are established from the source to the target layer and lateral connections within the target layer. Our work is inspired by “Case 1: STDP and rewiring operating simultaneously in the presence of correlated input” of [7]. The process of creating a topographic mapping is driven by correlated input from the devices in the source layer which generate spikes with Poisson-distributed inter-spike intervals according to spatially modulated firing rates following a Gaussian decay from a stimulus center that changes location randomly in regular intervals as illustrated in figure 4B. After network initialization, both feed-forward and lateral connections are continuously updated with synaptic plasticity (i.e., weight adjustments according to an STDP rule) and structural plasticity (i.e., changes to the connectivity itself as described in section 2.1.3 in line with the “fast” rewiring protocol of [7]). Starting from a rough topographic map, simulating the model leads to a refinement of receptive fields: each neuron’s preferred location approximates its ideal location. Note that the formation rule assumes wider feed-forward than lateral connections as shown in figure 4C and D. We deviate from previous implementations [22, 7] in the following points:
@@
  • Instead of nearest-neighbor spike pairing, we use all-to-all pairing of pre- and postsynaptic spikes to compute the weight updates of the STDP model [2].
  • We also use the formation rule to initialize the connectivity, but instead of fixing the initial in-degree, we apply the “pairwise Bernoulli” rule [1] using probabilities computed according to the formation probability given in table 1. This approach leads to on average $K=2\pi p_{\text{form}}\sigma_{\text{form}}$ synapses per neuron instead of $32$ and excludes multapses [1] from the beginning as these are not supported by the structural plasticity rule.
@@
  • We use a simulation time step of $h=0.1\,\mathrm{ms}$ (same as [22]) in contrast to $1\,\mathrm{ms}$ [7] and solve the neuron dynamics with the exponential Euler method. Our delay also corresponds to one time step.
@@ 3 Results
3.1 DVS gesture classification with DEEP R
Figure 5: DVS gesture classification results. (A) Accuracy. Bar heights show mean accuracy and error bars standard deviations, all calculated over five repeats with different seeds. The dashed line shows the mean test accuracy of the baseline dense network. (B) Training time. The sparse models shown have the same sparsity in $W^{\textrm{in}}$and $W^{\textrm{rec}}$.
@@ 3.1 DVS gesture classification with DEEP R
Figure 5A shows the training and testing accuracy of sparse and dense classifier networks (section 2.2.1) trained on the DVS gesture dataset [21] using mlGeNN (see section 2.1.1) with the GPU-parallelised DEEP R implementation described in section 2.1.2. We trained all models for $100$ epochs using a batch size of $512$ and the Adam optimizer [35] with default parameters. We also used the default e-prop parameters for all networks but, as the dataset does not provide a validation split, we employ a cross-validation methodology – where we leave out two users in each fold – to optimize the L1 regularization strength (after [16]).
@@
Except for the configuration with $5\,\mathrm{\%}$ input connectivity and $1\,\mathrm{\%}$ recurrent connectivity, using DEEP R significantly improves the performance of all sparse models trained on DVS gesture compared to the versions trained with fixed connectivity ($p<0.05$, paired, one-sided Wilcoxon signed rank test [38], $n=5$). Furthermore, as long as there are sufficiently many input connections ($\geq$$5\,\mathrm{\%}$ input connectivity), sparse models trained with DEEP R match the performance of the dense reference networks. Specifically, our model with $5\,\mathrm{\%}$ input connectivity and $1\,\mathrm{\%}$ recurrent connectivity obtains a test accuracy of $88.86\pm 0.3\,\mathrm{\%}$ compared to $88.94\pm 0.6\,\mathrm{\%}$ for the dense model. This is higher than the $88.0\,\mathrm{\%}$ reported by [34] for an event-based Gated Recurrent Unit (GRU) model and is achieved with around $90\times$ fewer parameters ($60\times 10^{3}$ vs. $5.5\times 10^{6}$).
@@
Figure 6A shows the training dynamics of three types of models and demonstrates that the training of all types is stable, although, as the literature predicts, the sparser models train more slowly [39]. Figure 6B shows how the rate of DEEP R rewiring changes over time with drastic changes in connectivity happening during the first $100$ batches but connectivity remaining more or less unchanged towards the end of training, matching the intuition that DEEP R is sampling network connectivity and then settling into a possible solution.
@@
We then measured computational speed. As shown in figure 5B, training models with $5\,\mathrm{\%}$ connectivity using e-prop and GeNN is up to $8\times$ faster than training a dense model and training models with $1\,\mathrm{\%}$ connectivity is up to $15\times$ faster. While the time spent running the DEEP R updates described in section 2.1.2 takes a maximum of $0.5\,\mathrm{\%}$ of the total training time, training the models with DEEP R is slightly slower because we (conservatively) double the maximum row length in the ragged sparse matrix data structure (see section 2.1) to allow space for rewiring.
3.2 Topographic map formation
@@
Figures 7 and 8 show the evolution of feed-forward and lateral connections, respectively. Following the rough topographic map initializion according to the formation probability given in table 1, the elimination rate gradually increases and then stabilizes, while the formation rate is mostly stable over the time course of the simulation, see panels A and B. The increased elimination probabilities in our implementation compared to [22] result in similar elimination and formation rates after the initial warm-up phase without the need for fixing the maximum in-degree. These rates are higher for the feed-forward than for the lateral connections due to the higher standard deviation of the Gaussian-shaped formation probability. In particular in the beginning of the simulation, it can be observed that connections established across larger distances get predominantly eliminated because the synaptic plasticity weakens these synapses triggered by the spatially correlated Poisson input. Note that the upper displays of panels A and B only count the changes in connectivity at certain spatial separations of the respective grid locations without taking into account the number of possible connections per distance. For instance, there is only one grid position at zero distance (for lateral connections this corresponds to autapses [1]) and four grid positions at a distance of one. Panel C shows how the connections per neuron evolve on average. The results for in-degrees and out-degrees are similar. The starting point matches the theoretical expectation, and it is the same for feed-forward and lateral because in both cases the parameters are chosen such that $p_{\text{form}}\sigma_{\text{form}}=1$. From this starting point, the number of feed-forward connections per neuron grows until saturation. In contrast, the number of lateral connections decreases after an initial increase until reaching a stable value. While figure 4C and D show connection probabilities and weights averaged over displacements for each grid position just for the initial and final state, panel D of figures 7 and 8 illustrate the temporal evolution across all y-displacements with zero x-displacement. The gradual refinement of the receptive fields is visible mostly close to the ideal location. In more distant places new synapses get formed, but their weights decrease from their high initial value due to the normal synaptic weight plasticity, and they get eventually eliminated again.

### v2

@@ 2.1.3 Synaptic rewiring for topographic maps
Figure 3: Illustration of data structures used for one custom connectivity update according to the synaptic rewiring model for topographic maps. Each panel shows a network sketch, the ragged matrix (number of presynaptic neurons $\times$ rmaximum out-degree), the bitfield (number of presynaptic neurons $\times$ number of postsynaptic neurons), and pseudocode. (A) Initial state. Five presynaptic neurons, sources $s_{j}$ with $0\leq j<5,$ are connected to four postsynaptic neurons, targets $t_{i}$ with $0\leq i<4$, using sparse connectivity with a maximum out-degree of three. Weights below the threshold $g_{\theta}$ are indicated in light red, others in light blue. A synapse slot from $s_{j}$ to $t_{i}$ is here referred to as $\left(s_{j},t_{i}\right)$. The bitfield is cleared. (B) Host assigns six rewiring attempts randomly to three presynaptic neurons ($s_{0}$, $s_{2}$, $s_{4}$). The presynaptic neurons then in parallel select their postsynaptic targets randomly and set the corresponding bits as candidates for rewiring attempts in the bitfield. Four existing synapses are selected for elimination (green; $\left(s_{0},t_{0}\right)$, $\left(s_{0},t_{2}\right)$, $\left(s_{4},t_{2}\right)$, $\left(s_{4},t_{3}\right)$) and two empty synapse slots for formation (yellow; $\left(s_{2},t_{2}\right),\left(s_{4},t_{0}\right)$). (C) Elimination attempts. $\left(s_{0},t_{0}\right)$ has a low elimination probability because of the strong weight (light blue) and remains, the three others have a higher elimination probability because of the lower weights (light red); $\left(s_{0},t_{2}\right)$ gets eliminated reducing the out-degree of $s_{0}$ in the ragged matrix from two to one; $\left(s_{4},t_{2}\right)$ gets eliminated, and the emerging gap in the row of $s_{4}$ is filled by inserting the last element $t_{3}$; $\left(s_{4},t_{3}\right)$ remains. The elimination bits get cleared in the bitfield. (D) Formation attempts. Assuming a spatial proximity of $s_{2}$ and $t_{2}$ and therefore a high formation probability, a new synapse (purple) is formed at $\left(s_{2},t_{2}\right)$, increasing the out-degree of $s_{2}$ from zero to one, but not at $\left(s_{4},t_{0}\right)$ because $s_{4}$ and $t_{0}$ are farther apart. The formation bits get cleared in the bitfield.
@@
We develop a GeNN-compatible variation of the synaptic rewiring model proposed by [undefaa], drawing inspiration from an implementation on the SpiNNaker platform [undefj]. This stochastic model combines a distance-dependent formation rule with a weight-dependent removal rule. The simulation loop performs custom connectivity updates at regular intervals according to a rewiring frequency. Each update triggers a fixed number of rewiring attempts between neurons from the presynaptic and postsynaptic population. While previous implementations [undefaa, undefj] handle each rewiring attempt separately, our algorithm aims to increase efficiency by parallelizing multiple attempts. Figure 3 illustrates the hybrid CPU (host) and GPU (device) algorithm processing these attempts and highlights the data structures. First, serial host code (host_update_code) uniformly distributes the attempts across all neurons in the presynaptic population. As we sample with replacement, multiple attempts per presynaptic neuron are possible. In contrast to earlier implementations [undefaa, undefj], our version does not impose a fixed maximum in-degree. GeNN represents connectivity in 2D-array structures; the rows correspond to presynaptic neurons and can be operated on independently and hence in parallel. Here, we introduce a bitfield as an auxiliary structure in a similar format to mark synaptic slots between specific pre- and postsynaptic neurons for wiring attempts. On the GPU (row_update_code), each presynaptic neuron accesses a unique section of the bitfield with as many bits as there are neurons in the postsynaptic population. The bits therefore encode synapse slots for potential synapses between the presynaptic neuron of the row considered and all postsynaptic neurons; each bit refers to a postsynaptic neuron index. The bitfield section is initialized to zero and then as many bits are set randomly according to a uniform distribution without replacement as there are wiring attempts assigned. Each synaptic slot is therefore considered at most once, unlike the earlier versions [undefaa, undefj] where each rewiring attempt is completely independent of the others.
@@
When the presynaptic neuron has set its bitfield row, the elimination rule is applied: A loop over all existing synapses of that presynaptic neuron identifies those that are set in the bitfield section, removes them dependent on an elimination probability compared to a random number and clears the respective bit. Weak synapses are more likely to be removed: the elimination probability is higher for depressed synapses with weights below the threshold than for strong, potentiated synapses. After all elimination attempts, the remaining set bits in the bitfield row point to synapse slots that were initially empty. Applying the formation rule, we then iterate over the whole bitfield row: If a bit is still set, we attempt to establish a synapse according to a Gaussian probability depending on the distance between the neurons and also clear that bit. The shorter the lateral distance between pre- and postsynaptic neuron, the more likely is a new connection. Our interpretation of the rewiring model does not enable the formation of multapses (i.e., multiple synapses per neuron pair [undef]). If a synapse exists, only the elimination rule is applied. To sum up, each custom connectivity update initiates a fixed number of rewiring attempts between pre- and postsynaptic neuron populations. The outcome of each attempt is a newly formed connection, a removed connection, or no change to the connectivity.
@@ 2.2.1 Recurrent classifier
Our previous work [undeft, undefu, undefv] showed that recurrent SNN classifiers can be trained on datasets such as the Spiking Heidelberg Digits [undefal] and DVS gesture [undefz].
Here we train SNN classifiers with a single recurrently connected hidden layers of 512 neurons on the N-MNIST and DVS gesture datasets. Following the approach outlined by [undefam], we spatially downsample the DVS gesture examples from $128\times 128$ to $32\times 32$ resolution and temporally downsample so that there is only one spike in each $1\,\mathrm{ms}$ simulation timestep. When training on DVS gesture, we use adaptive LIF (ALIF) neurons, the membrane voltage ($v$), adaptation variable ($a$) and spiking output ($z$) of which can be described by:
@@
where $\alpha=e^{-\frac{\textrm{dt}}{\tau_{\textrm{mem}}}}$represents the decay of the membrane with time constant $\tau_{\textrm{mem}}=20\,\mathrm{ms}$, $W^{\textrm{rec}}$ are the recurrent weights, $W^{\textrm{in}}$ are the weights connecting the input to the recurrent layer, $x^{t}$ are input spikes from the dataset, $v_{\textrm{thr}}=0.6$ is the baseline spiking threshold, $\rho=e^{-\frac{\textrm{dt}}{\tau_{\textrm{adapt}}}}$represents the adaptation variable decay with time constant $\tau_{\textrm{adapt}}=2000\,\mathrm{ms}$ and $\beta=0.0174$ is the adaptation strength. When neurons spike, $v_{\textrm{thr}}$ is subtracted from their membrane voltage $v$. In our classifiers, $W^{\textrm{rec}}$ and $W^{\textrm{in}}$ are sparse and initialised with a fixed probability of connection between each pair of pre and postsynaptic neurons e.g. we use $5\,\mathrm{\%}$ connectivity to indicate that there is a $5\,\mathrm{\%}$ chance of a pair of pre and postsynaptic neurons being connected. When training on the simpler N-MNIST task, we use standard LIF neurons, the dynamics of which can be obtained by setting $\beta=0$ and $\rho=0$.
@@
We calculate weight gradients $\Delta W^{\textrm{rec}}$ and $\Delta W^{\textrm{in}}$ using eligibility traces $\epsilon$ and $e$ calculated using e-prop [undefw] (shown here for $\Delta W^{\textrm{rec}}$) :
@@
where $\psi_{j}^{t}=\frac{0.3}{v_{\textrm{thr}}}\text{max}\left(0,1-\left|\frac{v_{j}^{t}-(v_{\textrm{thr}}+\beta a_{j}^{t})}{v_{\textrm{thr}}}\right|\right)$ is a surrogate gradient function, $\bar{z}^{t}$ is a version of $z^{t}$ low-pass filtered with a time constant of $\tau_{\textrm{mem}}$, $\eta$ is a learning rate, $B=\left(W^{\textrm{out}}\right)^{T}$, $\pi_{k}^{t}=\textrm{softmax}(y_{1}^{t},\ldots,y_{K}^{t})$, $\pi^{*,t}$ are one-hot encoded labels and $\bar{e}^{t}$ is a version of $e^{t}$ low-pass filtered with a time constant of $\tau_{\textrm{mem}}$. When using LIF neurons, only a single simpler eligbility trace is required:
$e_{ji}^{t+1}$ $=\psi_{j}^{t+1}\bar{z}_{i}^{t}$ (8)
@@
All gradients are applied to the underlying variables using the Adam optimizer [undefan] and $\Delta W^{\textrm{rec}}$ and $\Delta W^{\textrm{in}}$provide the inputs to the DEEP R algorithm described in section 2.1.2.
@@ B Neuron, input and plasticity models
Table 1: Topographic map model description and parameters following the guidelines of [undefao].
@@ F Synaptic plasticity
$B$ $1.2$ Ratio of potentiation to depression es defined in [undefap]
@@ G Structural plasticity
Elimination probability for depressed synapses ($g_{\text{syn}}<g_{\theta}$); value from [undefaa] increased by factor
@@
Elimination probability for potentiated synapses ($g_{\text{syn}}\geq g_{\theta}$); value from [undefaa] increased by factor
@@ H Simulation
The network model illustrated in figure 4A and its parameters closely match the earlier accounts on topographic map formation of [undefaa, undefj]. In this section, we provide a summary and emphasize the differences, while tables 1 and 2 give a complete overview of the model description and parameters used in our work in the format frequently used in computational neuroscience [undefao]. The model consists of two square layers with nodes located at grid positions: model neurons in the target layer and spike-generating devices in the source layer. Feed-forward connections are established from the source to the target layer and lateral connections within the target layer. Our work is inspired by “Case 1: STDP and rewiring operating simultaneously in the presence of correlated input” of [undefj]. The process of creating a topographic mapping is driven by correlated input from the devices in the source layer which generate spikes with Poisson-distributed inter-spike intervals according to spatially modulated firing rates following a Gaussian decay from a stimulus center that changes location randomly in regular intervals as illustrated in figure 4B. After network initialization, both feed-forward and lateral connections are continuously updated with synaptic plasticity (i.e., weight adjustments according to an STDP rule) and structural plasticity (i.e., changes to the connectivity itself as described in section 2.1.3 in line with the “fast” rewiring protocol of [undefj]). Starting from a rough topographic map, simulating the model leads to a refinement of receptive fields: each neuron’s preferred location approximates its ideal location. Note that the formation rule assumes wider feed-forward than lateral connections as shown in figure 4C and D. We deviate from previous implementations [undefaa, undefj] in the following points:
@@
  • Instead of nearest-neighbor spike pairing, we use all-to-all pairing of pre- and postsynaptic spikes to compute the weight updates of the STDP model [undefa].
  • We also use the formation rule to initialize the connectivity, but instead of fixing the initial in-degree, we apply the “pairwise Bernoulli” rule [undef] using probabilities computed according to the formation probability given in table 1. This approach leads to on average $K=2\pi p_{\text{form}}\sigma_{\text{form}}$ synapses per neuron instead of $32$ and excludes multapses [undef] from the beginning as these are not supported by the structural plasticity rule.
@@
  • We use a simulation time step of $h=0.1\,\mathrm{ms}$ (same as [undefaa]) in contrast to $1\,\mathrm{ms}$ [undefj] and solve the neuron dynamics with the exponential Euler method. Our delay also corresponds to one time step.
@@ 3 Results
3.1 Classification with DEEP R
Figure 5: Classification accuracy. Bar heights show mean accuracy and error bars standard deviations, all calculated over five repeats with different seeds. The dashed line shows the mean test accuracy of the baseline dense network. (A) N-MNIST. (B) DVS-gesture.
@@ 3.1 Classification with DEEP R
Figure 7: Performance and weight savings of sparse DVS gesture training. Sparse models shown have the same sparsity in $W^{\textrm{in}}$and $W^{\textrm{rec}}$. (A) Training time.(B) The number of weights in models.
@@
Figure 5 shows the training and testing accuracy of sparse and dense classifier networks (section 2.2.1) trained on the N-MNIST [undefy] and DVS gesture datasest [undefz] using mlGeNN (see section 2.1.1) with the GPU-parallelised DEEP R implementation described in section 2.1.2. We trained all models for $100$ epochs using a batch size of $512$ and the Adam optimizer [undefan] with default parameters. We also used the default e-prop parameters for all networks only optimizing the L1 regularization strength on a validation set. Neither dataset contains a dedicated validation split so, as N-MNIST has a large number of examples, we simply held out 10000 random example whereas, for DVS gesture, we employed a cross-validation methodology – where we leave out two users in each fold (after [undeft]).
@@
Except for the model trained on DVS gesture with $5\,\mathrm{\%}$ input connectivity and $1\,\mathrm{\%}$ recurrent connectivity; and the model trained on N-MNIST with $5\,\mathrm{\%}$ input connectivity and $5\,\mathrm{\%}$ recurrent connectivity, using DEEP R significantly improves the performance of all sparse models compared to the versions trained with fixed connectivity ($p<0.05$, paired, one-sided Wilcoxon signed rank test [undefaq], $n=5$). Furthermore, as long as there are sufficiently many input connections ($\geq$$5\,\mathrm{\%}$ input connectivity), sparse models trained with DEEP R matched the performance of the dense reference networks on both datasets. Specifically, our models with $5\,\mathrm{\%}$ input connectivity and $1\,\mathrm{\%}$ recurrent connectivity trained on DVS gesture obtains a test accuracy of $88.86\pm 0.3\,\mathrm{\%}$ compared to $88.94\pm 0.6\,\mathrm{\%}$ for the dense model. This is higher than the $88.0\,\mathrm{\%}$ reported by [undefam] for an event-based Gated Recurrent Unit (GRU) model and is achieved with around $90\times$ fewer parameters ($60\times 10^{3}$ vs. $5.5\times 10^{6}$).
@@
Figure 6A shows the training dynamics of three types of models on the DVS gesture dataset and demonstrates that the training of all types is stable, although, as the literature predicts, the sparser models train more slowly [undefar]. Figure 6B shows how the rate of DEEP R rewiring changes over time with drastic changes in connectivity happening during the first $100$ batches but connectivity remaining more or less unchanged towards the end of training, matching the intuition that DEEP R is sampling network connectivity and then settling into a possible solution.
@@
We then measured computational speed of DVS gesture training. As shown in figure 7A, training models with $5\,\mathrm{\%}$ connectivity using e-prop and GeNN is up to $8\times$ faster than training a dense model and training models with $1\,\mathrm{\%}$ connectivity is up to $15\times$ faster. While the time spent running the DEEP R updates described in section 2.1.2 takes a maximum of $0.5\,\mathrm{\%}$ of the total training time, training the models with DEEP R is slightly slower because we (conservatively) double the maximum out-degree in the ragged sparse matrix data structure (see section 2.1) to allow space for rewiring. Figure 7B shows the number of weights in each model configuration. Due to the padding, the number of weights and hence memory savings are not directly proportional to the sparsity. Nonetheless, DEEP R models with $1\,\mathrm{\%}$ connectivity still require $11\times$ fewer weights than if they were implemented using dense matrices and those with $5\,\mathrm{\%}$ connectivity $4\times$ fewer.
Overall, in GeNN, connection sparsity significantly reduces training times and memory usage. This represents a significant advantage compared to SNN simulation libraries such as SNNtorch [undefae], NORSE [undefaf] or JAXsnn [undefag] which are built on standard Machine Learning toolkits like PyTorch or JAX and thus only provide efficient support for dense weight matrices.
@@
3.2 Topographic map formation
Figures 8 and 9 show the evolution of feed-forward and lateral connections, respectively. Following the rough topographic map initializion according to the formation probability given in table 1, the elimination rate gradually increases and then stabilizes, while the formation rate is mostly stable over the time course of the simulation, see panels A and B. The increased elimination probabilities in our implementation compared to [undefaa] result in similar elimination and formation rates after the initial warm-up phase without the need for fixing the maximum in-degree. These rates are higher for the feed-forward than for the lateral connections due to the higher standard deviation of the Gaussian-shaped formation probability. In particular in the beginning of the simulation, it can be observed that connections established across larger distances get predominantly eliminated because the synaptic plasticity weakens these synapses triggered by the spatially correlated Poisson input. Note that the upper displays of panels A and B only count the changes in connectivity at certain spatial separations of the respective grid locations without taking into account the number of possible connections per distance. For instance, there is only one grid position at zero distance (for lateral connections this corresponds to autapses [undef]) and four grid positions at a distance of one. Panel C shows how the connections per neuron evolve on average. The results for in-degrees and out-degrees are similar. The starting point matches the theoretical expectation, and it is the same for feed-forward and lateral because in both cases the parameters are chosen such that $p_{\text{form}}\sigma_{\text{form}}=1$. From this starting point, the number of feed-forward connections per neuron grows until saturation. In contrast, the number of lateral connections decreases after an initial increase until reaching a stable value. While figure 4C and D show connection probabilities and weights averaged over displacements for each grid position just for the initial and final state, panel D of figures 8 and 9 illustrate the temporal evolution across all y-displacements with zero x-displacement. The gradual refinement of the receptive fields is visible mostly close to the ideal location. In more distant places new synapses get formed, but their weights decrease from their high initial value due to the normal synaptic weight plasticity, and they get eventually eliminated again.